OpenStudy (anonymous):

Which pair of lines is parallel? y = 4x + 1 and y = 4x =5 y = -2x and 2y - 2x = -2 y = 1/4x + 2 and y -2= 1/2x y = 1/5x + 1 and 5y _ x = 10

OpenStudy (anonymous):

You can tell that 2 lines are parallel if their slopes have the same magnitude

OpenStudy (anonymous):

first one

OpenStudy (texaschic101):

parallel lines will have the same slope, so just find the slopes of these equations and see which ones are the same

OpenStudy (anonymous):

How do you find the slope of equations?

OpenStudy (texaschic101):

P13...you need to stop just giving answers without explaining

OpenStudy (texaschic101):

if the equations are in y = mx + b form, the m position is the slope for example : y = 4x + 1 the number in the m position is 4, so the slope is 4 if it is not in y = mx + b form, you need to put it into y = mx + b form for example : y - 2 = 1/2x -- add 2 to both sides y = 1/2x + 2 ...the slope here is 1/2 does that make sense to you ?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (texaschic101):

no problem...glad to help :)